# A progress on the binary Goldbach conjecture

### v6

@@ 4. $\bullet$
then there exists an axis $\mathbb{L}_{[v_{2}],[w_{1}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})$ with $0<s<t$. Hence also $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})$ is not empty.
@@
By virtue of the requirements $w_{1},v_{2}\in\mathbb{C}_{\mathbb{B}}$ and $n+s=\Re(w_{1})+\Re(v_{2})$ there is an axis $\mathbb{L}_{[v_{2}],[w_{1}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})$ and hence holds $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})\neq\emptyset$. And from $\mathbb{B}\subset\mathbb{M}$ follows immediately $\mathbb{C}_{\mathbb{B}}\subset\mathbb{C}_{\mathbb{M}}$ and therefore holds also $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})\neq\emptyset$. This completes the proof. ∎
@@
then there exists an axis $\mathbb{L}_{[w_{1}],[z_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{P}})$ with $0<s<t$.
@@ 4. The generalized squeeze principle and applications
then there exists an axis $\mathbb{L}_{[v_{2}],[w_{1}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})$ with $0<s<t$. Hence $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})$ is also not empty.

### v7

@@ 4. $\bullet$
Remark 3.1.
Let $\mathcal{C}(n,\mathbb{B})$ and $\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{B}})$ be a CoP and a corresponding cCoP, respectively. It is clear that $\mathcal{C}(n,\mathbb{B})\neq\emptyset$ if and only if $\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{B}})\neq\emptyset.$ To see this, observe that if $\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{B})$, then $\mathbb{L}_{[x+it],[y-it]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{B}})$. Conversely if $\mathbb{L}_{[z_{1}],[z_{2}]}~{}\hat{\in}~{}\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{B}})$ then $z_{1}+z_{2}=n$ implying that $\Im(z_{1})=-\Im(z_{2})$ and so $\Re(z_{1})+\Re(z_{2})=n$. This immediately implies that there is an axes $\mathbb{L}_{[\Re(z_{1})],[\Re(z_{2})]}\in\mathcal{C}(n,\mathbb{B})$. This equivalence will be subtly employed in the rest of the paper.
@@
then there exists an axis $\mathbb{L}_{[\Re(v_{2})],[\Re(w_{1})]}~{}\hat{\in}~{}\mathcal{C}(n+s,\mathbb{B})$ with $0<s<t$. Hence the cCoP $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})$ not empty and so is the cCoP $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})$.
@@
By virtue of the requirements $w_{1},v_{2}\in\mathbb{C}_{\mathbb{B}}$ and $n+s=\Re(w_{1})+\Re(v_{2})$ there is an axis $\mathbb{L}_{[\Re(v_{2})],[\Re(w_{1})]}~{}\hat{\in}~{}\mathcal{C}(n+s,\mathbb{B})$ and hence holds $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{B}})\neq\emptyset$. Since $\mathbb{B}\subset\mathbb{M}$, it follows immediately that $\mathbb{C}_{\mathbb{B}}\subset\mathbb{C}_{\mathbb{M}}$ and therefore $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})\neq\emptyset$. This completes the proof. ∎
@@
then there exists an axis $\mathbb{L}_{[\Re(w_{1})],[\Re(z_{2})]}~{}\hat{\in}~{}\mathcal{C}(n+s,\mathbb{P})$ with $0<s<t$.
@@ 4. The generalized squeeze principle and applications
then there exists an axis $\mathbb{L}_{[\Re(v_{2})],\Re([w_{1})]}~{}\hat{\in}~{}\mathcal{C}(n+s,\mathbb{B})$ with $0<s<t$. Hence $\mathcal{C}^{o}(n+s,\mathbb{C}_{\mathbb{M}})$ is also not empty.